Prev Next

SW&IT Meeting Minutes 2021-06-07

Attendees: David, Matthieu, Michael, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210531
LR20210531 - Bec, Matthieu 40.0
LR20210531 - Edmeades, David 0.0
LR20210531 - Grenz, Paul 20.0
LR20210531 - Hooper, Stephen 0.0
LR20210531 - Zhang, Yang 0.0
LR20210531 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • dms.tucson on KVM
  • robs disk usage and ~/.luci auto-cleaning

Team reports

MB

  • Last week
    • OC, Management
    • cobbler, jenkins, docker infra10
    • lit2.dev, facsum
    • s/w interviews
    • investigate some AO/TCS issues?
  • This week
    • lit2.dev, facsum
    • ATT task lists, meeting
    • infra10/ovirt follow up

YZ

  • Previous:
    • Vacation
  • Next:
    • Continuing previous TMS work (bino-mode software, mirror movement verification).
    • Continuing some MODS Fedora/robs testing and fixing (in discussion with Olga).
    • Pickup some OSCO work.

XZ

  • Last week
    • Analysis the AO/AdSec issues from May 24-31
    • Document the AO/AdSec events from May 24-31
    • AO support
  • This week
    • Define DX lab test criterias
    • Analysis SX/DX data to define DX lab test criterias
    • AO support/meetings

PSG

  • Last week:
    • Tested swing arms with Gustavo.
    • Motion now works,
    • Swing arms can get into a state where it does not recognize stow pins.
    • This is not something I can fix with software.
  • This week:
    • Installing and testing full OSS on tcs-test.

DE

  • Last week
    • Office trip
      • Drive swap
      • Network work
        • Unknown issue with ex-538 switch, pulled from service for UITS to look at. Reply from UITS contact--he is going to look at it where it is in 580.
        • Replacement 538 switch no problems, still working--checked since SH urged caution about premature jubilation
        • MW office switch install, cable untangling. South side working, but he has gear on the north that we'll probably just pull cables to. Also possible followup on iMac IP config.
        • DC office, unable to replicate problem, DC verified connectivity later. Registered all MAC addresses he has
    • Opened conversation with Ketiv re: Vault licensing. Major turnover there, have to train new guy up.
    • Power bump on mountain
      • Equipment damage: Eliott desktop and L10 switch
      • Eliott will drop off workstation for us to look at.
  • This week
    • Switch deadline looming, so I'll plan another office trip for my office as well as poke around to see if there are any other workgroup switches I can swap out
    • Ketiv followup
    • Veeam/Backblaze checkup
    • MarkS startup
    • Robert handholding
    • Reminder: this week is an all-hands Friday

SH

  • Late back from vacation, so a short week:
    • Thursday - broken switch sw1001, ENV stuff (with Josh the night before as well), ovirt, sorting out DNS.
    • Friday - interviewing SWG candidates, no chance on getting ovirt tweaked, not much with sw1001 except made an appointment with the mountain guys for Sunday.
    • Saturday - Jay, and Tim replaced the sw1001 switch once I configured the new to us one.

next:

  • ovirt -- get the engine fired up somehow.
  • DIMM
  • LBC cabinet switches -- maybe (for SSD)
  • Elliot's computers stuff, dms KVM while I am in the office, LeRoy and weather.
Topic revision: r8 - 07 Jun 2021, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback